OpenStudy (calculusxy):

Find coordinates for a point that is three times as far from the origin as (2, 3) is. Describe the configuration of all such points.

OpenStudy (calculusxy):

@phi

OpenStudy (phi):

do you know how far away (2,3) is from the origin ? (I would use the "distance formula")

OpenStudy (calculusxy):

\(\sqrt{13}\)

OpenStudy (phi):

and three times that distance is \[ 3 \sqrt{13} \] they want a point that far from the origin I can think of a few "easy" points. For example, make the y value 0, what x value should we use ?

OpenStudy (calculusxy):

\((0, 3\sqrt{13})\)

OpenStudy (calculusxy):

Or \((\pm 3\sqrt{13}, 0)\)

OpenStudy (calculusxy):

Let me make the \((0, 3\sqrt{13})\) into \((0, \pm 3\sqrt{13})\)

OpenStudy (calculusxy):

We could make a circle right by connecting all four of those points?

OpenStudy (phi):

ok, that is on the y-axis, but it works also and they also want to know what the "locus of points" i.e. all points equidistant from a common point (also known as the "center")

OpenStudy (calculusxy):

The radii would all be \(3\sqrt{13}\)

OpenStudy (phi):

yes, a circle is made up of points equally far from the center (the origin in this case)

OpenStudy (calculusxy):

So would be have something like \(x^2 + y^2 = (3\sqrt{13})^2\)?

OpenStudy (phi):

you could, but they don't ask for an equation (but giving the equation is ok)

OpenStudy (calculusxy):

Ok. So what do they want when they mean "describe the configuration"?

OpenStudy (phi):

I think they want: All points 3 times farther from the origin than the point (2,3) form a circle, with the center at (0,0), and radius 3 sqr(13)

OpenStudy (calculusxy):

Yes that makes so much sense :) Thank you!